# Covers of surfaces

### v1

@@
Abstract
We study the homeomorphism types of certain covers of (always orientable) surfaces, usually of infinite-type. We show that every surface with non-abelian fundamental group is covered by every noncompact surface, we identify the universal abelian covers and the $\mathbb{Z}/n\mathbb{Z}$-homology covers of surfaces, and we show that non-locally finite characteristic covers of surfaces have four possible homeomorphism types.
1 Introduction
We are interested in the question of when a given surface covers another given surface, either in general or with constraints on the type of covering. Our first theorem is as follows.
Theorem 1.1 (Everything covers everything).
@@ 1 Introduction
A borderless surface is a topological $2$-manifold with empty boundary. We stick with orientable surfaces for simplicity. The only orientable surfaces $S$ not covered by the theorem are the disk, the annulus, and the torus, each of which is only covered by itself and the former surface.
If $S$ is any surface, there are certain natural covers coming from homology. The universal abelian cover of $S$ is the cover $S^{ab}\to S$ corresponding to the commutator subgroup of $\pi_{1}(S)$. We prove:
@@
  4. 4.
    if $S$ is a finite-type surface with one puncture and genus at least $2$, then $S^{ab}$ is the spotted Loch Ness monster surface,
  5. 5.
@@
Note that the commutator subgroup of $\pi_{1}(S)$ is the kernel of the map $\pi_{1}(S)\to H_{1}(S,\mathbb{Z}),$ so the universal abelian cover could also be called the “$\mathbb{Z}$-homology cover” of $S$. We also consider the $\mathbb{Z}/n\mathbb{Z}$-homology cover $\tilde{S}\to S$ corresponding to the kernel of the map
@@
When $S$ has finite-type, the $\mathbb{Z}/n\mathbb{Z}$-homology cover $\tilde{S}\to S$ is a finite cover, and one can figure out the topological type of $\tilde{S}$ via an Euler characteristic argument and by analyzing how boundary components lift. On the other hand, when $\tilde{S}$ has infinite-type, its $\mathbb{Z}/n\mathbb{Z}$-homology covers all have infinite degree, and it turns out their topology is quite constrained.
@@
Suppose that a cover $\tilde{S}\to S$ corresponds to the subgroup $N<\pi_{1}(S,p)$. We say that the cover is characteristic if $N$ is characteristic, i.e., if $N$ is invariant under every automorphism of $\pi_{1}(S,p)$. The $\mathbb{Z}$ and $\mathbb{Z}/n\mathbb{Z}$-homology covers above are all characteristic. Motivated by Theorems 1.2 and 1.3, we ask:
@@
Above, an automorphism of $\pi_{1}(S,p)$ is called geometric if it is induced by a homeomorphism $S\to S$ that preserves the basepoint $p$. We call a subgroup $N<\pi_{1}(S,p)$ geometrically characteristic if $N$ is invariant under all geometric automorphisms of $\pi_{1}(S,p)$. So, ‘geometrically characteristic’ is a weaker condition than ‘characteristic’. A cover $\tilde{S}\to S$ is geometrically characteristic if it corresponds to a geometrically characterstic subgroup of $\pi_{1}(S,p)$.
It follows quickly from Theorem 1.5 that any non-universal characteristic cover $\tilde{S}$ of a closed surface is homeomorphic to the Loch Ness monster surface. This was proved earlier by Atarihuana–García–Hidalgo–Quispe–Maluendas [2, Prop 6.1], using the fact that $\tilde{S}$ is also a characteristic cover of a triangle orbifold. However, their argument does not extend to the finite-type setting.
@@
Our interest in Theorem 1.7 is partly inspired by work on big mapping class groups of surfaces. In particular, Aramayona–Leininger–McLeay [1, Proposition 3.3] show that the $\mathbb{Z}/2\mathbb{Z}$ homology cover of the Loch Ness monster surface is the Loch Ness monster surface. They use this to show that the mapping class group $G$ of the once-punctured Loch Ness monster surface admits a continuous injective homomorphism $G\to G$ that is not surjective. We note that there is a gap in the published version of their proof of the covering statement, but they were able to fix it quickly when we notified them about it. Alternatively, their Proposition 3.3 follows from Theorem 1.3, or even from Theorem 1.5 or [2, Proposition 6.1], since the Loch Ness monster surface is itself a characteristic cover of a genus two surface.
Surprisingly, to our knowledge there are not many previous results in which the homeomorphism types of infinite covers of surfaces are studied. In particular, we were surprised that a result like Theorem 1.1 does not currently exist in the literature. However, there are a few results out there that are related to ours. First, a classification of the possible homeomorphism types of infinite-degree regular covers of a closed surface is readily deduced from the fact that the cover and the deck group are quasi-isometric (for instance, see [4, Proposition 5.2]). Then there are older papers of Maskit [9] and Papakyriakopoulos [10] on identifying when certain covers of surfaces are planar; their motivation was some interesting applications to $3$-manifolds, including a new proof of the Loop Theorem. There are the papers of Aramayona–Leininger–McLeay [1] and Atarihuana et al [2] referenced above, and there is also a forthcoming paper of Ghaswala–McLeay in which (among other things) the authors construct finite-degree covers from certain maps between potential end spaces of surfaces.
1.1 Ideas of the proofs
@@
Theorem 1.1, in which we show that every noncompact surface covers every surface with non-abelian fundamental group, is very hands-on. Every surface $S$ with non-abelian fundamental group has an essential pair of pants and so is covered by an open pair of pants. The universal abelian cover of an open pair of pants is the Loch Ness monster surface, which is covered by the blooming Cantor tree surface, and every noncompact surface $\tilde{S}$ can be $\pi_{1}$-injectively embedded in the blooming Cantor tree surface, and hence covers the blooming Cantor tree surface. (Here the embedding is as a component of the complement of a set of closed curves and bi-infinite arcs.) Composing, we get a covering map $\tilde{S}\to S$.
Theorem 1.2, in which we identify the universal abelian covers $S^{ab}$ of surfaces $S$, is proved as follows. When $S$ has finite-type, one can identify $S^{ab}$ using the fact that the deck group is free abelian, along with the knowledge of whether the peripheral curves of $S$ lift or not. We deal with infinite-type $S$ by exhausting them with finite-type subsurfaces and saying that $S^{ab}$ is exhausted by the universal abelian covers of the subsurfaces.
@@
Say for simplicity that $S$ is closed. We want to prove that $\tilde{S}$ has one end. Here, $\tilde{S}$ is quasiisometric to the deck group $Q$. If $Q$ has more than one end, it admits an action on a tree $T$ with finite edge stabilizers and no global fixed point, by Stallings’s Theorem. We then construct a $Q$-equivariant map $\tilde{S}\to T$ that is transverse to the union of edge midpoints, and let $\tilde{C}\subset\tilde{S}$ be the preimage of the set of edge midpoints, which is then a properly embedded $1$-submanifold invariant under $Q$. Since edge stabilizers are finite, each component of $\tilde{C}$ is a simple closed curve. Using similar methods and an inductive argument, we can enlarge $\tilde{C}$ until every component of $\tilde{S}\smallsetminus\tilde{C}$ is one-ended.
@@
Theorem 1.7, in which we identify non-locally finite characteristic covers of infinite-type surfaces, is proved by exhausting by finite-type surfaces and using Theorem 1.5, in a way analogous to the infinite-type case of Theorem 1.2.
1.2 Outline of the Paper
In §2 we review the necessary background. We define the surfaces mentioned in the introduction, set conventions, and review actions on trees, ends, Stallings’s Theorem, and the classification of infinite-type surfaces. In §3 we study characteristic covers of surfaces, proving Theorems 1.5, 1.7, 1.2 and 1.3 in subsections 3.1, 3.2, 3.3 and 3.4, respectively. In §4 we prove Theorem 1.1, and then §A is our appendix on ends via group cohomology.
1.3 Acknowledgements
@@
2 Background
2.1 Conventions
All surfaces within this work are assumed to be connected, second countable, orientable, and possibly with boundary. We say the surface is borderless if it has no boundary, and remind the reader that some of the surfaces which appear in our proofs may have non-empty boundary with possibly non-compact boundary components. All borderless surfaces, whether finite or infinite type, are classified by their genus and the topology of their space of ends, see Theorem 2.2. We let $S^{b}_{g,p}$ denote the finite-type surface with $b$ boundary components, $g$ genus, and $p$ punctures.
@@ 2.1 Conventions
  3. 3.
    The spotted Loch Ness monster surface is obtained from the Loch Ness monster surface by removing an infinite, discrete set of points.
  4. 4.
@@
2.2 Ends of a topological space
@@
[5, Prop. 8.29] Given two proper geodesic metric spaces $X_{1}$ and $X_{2}$, then any quasi-isometry $f:X_{1}\to X_{2}$ induces a homeomorphism $f_{\mathcal{E}}:\mathcal{E}(X_{1})\to\mathcal{E}(X_{2})$.
2.3 Classification of surfaces
@@
Moreover, for every $n\in\{0,\ldots,\infty\}$ and for every compact, totally disconnected metrizable space $X$, and closed subset $C\subset X$, there is an orientable surface $S$ such that $g(S)=n$ and $(\mathcal{E}(S),\mathcal{E}_{g}(S))$ is homeomorphic to $(X,C)$.
@@ 2.4 Preliminaries on covers
2.5 Group actions on trees
@@
2.5.1 Stallings’s Theorem and its relative version
@@
This is really a result of Swarup [15], but Swarup’s result is stated in terms of group cohomology with group ring coefficients. In Appendix A, we give a gentle introduction to ends via group cohomology and show how the formulation above follows from Swarup’s result.
@@
3 Characteristic covers
In this section we study characteristic covers of surface, proving Theorems 1.5, 1.7, 1.2 and 1.3 in sections 3.1, 3.2, 3.3 and 3.4, respectively.
3.1 Finite-type surfaces
@@
Let $Q$ be the associated deck group of $\tilde{S}$. First assume that $\tilde{S}$ has no compact boundary components, so $\text{int}(\tilde{S})=\tilde{S}_{\circ}$. As $Q$ is infinite and acts cocompactly on $\tilde{S}$, the genus of $\text{int}(\tilde{S})$ is either 00 or infinity, so the classification of surfaces (Theorem 2.2) implies that it is either the disc or the Loch Ness monster surface. Now, suppose that $\tilde{S}$ has a compact boundary component. As $Q$ is infinite, $\tilde{S}$ has infinitely many compact boundary components. Gluing discs to each compact boundary component, we obtain a surface without boundary, call it $Y$. By Proposition 3.1, $Y$ has a single end, and again the classification of surfaces (Theorem 2.2) implies that $Y$ is either the disc or the Loch Ness monster surface. To finish, observe that $\text{int}(\tilde{S})$ is obtained from $Y$ by puncturing all the infinitely many discs, so then $\text{int}(\tilde{S})$ is either the flute surface or the spotted Loch Ness monster surface. ∎
@@
Let $\pi:\tilde{S}\to S$ be a regular cover of a compact surface $S$ with deck group $Q$. Suppose $Q$ acts on a tree $T$ with finite edge stabilizers, no edge inversions, no global fixed point, and such that the projection to $Q$ of every peripheral element of $\pi_{1}(S)$ acts elliptically on $T$. Then there exists a $Q$-equivariant map $\tilde{f}:\tilde{S}\to T$ such that each component of $\partial\tilde{S}$ maps to a vertex under $\tilde{f}$. Moreover, if $Z$ is the set of edge midpoints in $T$, then the preimage $\tilde{C}:=\tilde{f}^{-1}(Z)$ is a properly embedded $1$-submanifold of $\tilde{S}$, each component of which is a simple closed curve, and its projection $C:=\pi(\tilde{C})$ is a compact $1$-submanifold of $S$ each component of which is non-peripheral and essential.
Above, there may be components of $C$ that are parallel, which is why we call $C$ a $1$-submanifold and not a multicurve.
@@
Figure 1: The map $\tilde{f}_{1}$ from Proposition 3.3 restricted to a triangle in the triangulation $\tilde{\tau}$; the green, blue, and red regions map onto the green, blue, and red segments, respectively, in the tripod; the restriction to any one of the six subtriangles is an orthogonal projection map. The preimage of an edge midpoint $z\in T$ is shown in orange.
@@
First, suppose that there is a component of $C_{1}$ that bounds a disk. Let $\Delta_{1},\ldots,\Delta_{k}$ be the outermost disks bounded by components of $C_{1}$, so that if $c$ is a component of $C_{1}$ bounding a disk, then there exists a unique $j\in\{1,\ldots,k\}$ such that $c\subset\Delta_{j}$. Choose a lift $\tilde{\Delta}_{j}$ in $\tilde{S}$ for each of the $\Delta_{j}$, and for $q\in Q$, let $\tilde{\Delta}_{j}^{q}=q\cdot\tilde{\Delta}_{j}$. Let $e_{j}^{q}$ be the edge with vertices $v_{j}^{q},w_{j}^{q}\in T$ whose midpoint is $\tilde{f_{1}}(\tilde{\partial}\Delta_{j}^{q})$. By the construction of $C_{1}$, all the $1$-cells in $\tilde{\tau}$ that intersect ~$\tilde{\partial}\Delta_{j}^{q}$ are mapped to segments in $T$ that all contain the edge $e_{j}^{q}$. If the 00-cells in $\tilde{\Delta}_{j}^{q}$ are mapped into the $v_{j}^{q}$ side of $T\smallsetminus\{e_{j}^{q}\}$, we can define a new map
@@
Since the action of the deck group $Q$ produced by Lemma 3.2 has no global fixed point, the image of the map in Proposition 3.3 has to contain an edge midpoint, the $1$-submanifold $C\subset S$ it produces is non-empty. Take $\gamma$ to be any component of $C$. Any component of $\pi^{-1}(\gamma)$ is a component of $\tilde{C}$, and hence is compact. ∎
@@
If $\pi:\tilde{S}\to S$ is a regular cover of a compact surface and $C\subset S$ is a compact $1$-submanifold such that every component of $\pi^{-1}(C)\subset\tilde{S}$ is compact. Then there exists a compact $1$-submanifold $D\subset S$ containing $C$ such that
@@
3.1.1 Proof of Proposition 3.1
We argue by way of contradiction: assume that $\tilde{S}_{\circ}$ has more than one end. Let $Q$ be the deck group associated to the covering $\pi$. Apply Proposition 3.3 to obtain a tree $T$, a peripherally elliptic action of $Q$ on $T$ with no global fixed point, and the 1-submanifolds $\tilde{C}$ and $C$ of $\tilde{S}$ and $S$, respectively, and apply Lemma 3.5 to get the $1$-submanifold $D$ on $S$. Recall that the action of $Q$ on $T$ has finite edge stabilizers, no edge inversions, no global fixed point, and every element of $Q$ that is the image of a peripheral element in $\pi_{1}(S)$ acts elliptically on $T$. The contradiction will arise by showing that the action of $Q$ has a global fixed point.
@@
Figure 2: Standard generators for $\pi_{1}(S)$ for $S$ on a genus two surface with three punctures are given by the elements $a_{1},\ldots,a_{6}$.
Figure 3: A generating set for a thrice-punctured torus consisting of non-separating curves, namely $a_{1},a_{2},a_{3},a_{4}$, where products of pairs of distinct generators are either non-separating, or are freely homotopic to $\delta$. For a similar example on a twice-punctured torus, just delete one of the two boundary components that is not $\delta$.
@@
3.2 Non-locally finite covers of infinite-type surfaces

### v3

@@
Abstract.
We study the homeomorphism types of certain covers of (always orientable) surfaces, usually of infinite type. We show that every surface with non-abelian fundamental group is covered by every noncompact surface, we identify the universal abelian covers and the $\mathbb{Z}/n\mathbb{Z}$-homology covers of surfaces, and we show that non-locally finite characteristic covers of surfaces have four possible homeomorphism types.
1. Introduction
We are interested in the question of when a given surface covers another given surface, either in general or with constraints on the type of covering. As motivation, let us mention the following elementary result.
Proposition 1.1 (Everything covers everything).
@@ 1. Introduction
A borderless surface is a topological 222-manifold with empty boundary. We stick with orientable surfaces for simplicity. The only orientable surfaces $S$ not covered by the theorem are the disk, the annulus, and the torus, each of which is only covered by itself and the former surface.
Proposition 1.1 is not surprising, and it is not difficult. Related results already appear in the literature: Goldman [11] proves the result for closed $S$, which may also be non-orientable, and similar results are implicit in some papers on foliations, see e.g. [9], [1], [2], and [12]. With some extra work, the methods used by Goldman and others could likely yield our statement. However, our proof is short, constructive and the result motivates well the rest of our work.
In another direction, every surface $S$ has distinguished covers coming from especially important subgroups of $\pi_{1}(S)$. As a simple first example, the universal abelian cover of $S$ is the cover $S^{ab}\to S$ corresponding to the commutator subgroup of $\pi_{1}(S)$. In contrast to the flexibility displayed in Proposition 1.1, universal abelian covers only have five possible homeomorphism types!
@@
  4. (4)
    if $S$ is a finite-type surface with one puncture and genus at least 222, then $S^{ab}$ is the spotted Loch Ness monster surface,
  5. (5)
@@
The proof uses some basic geometric group theory to identify the universal abelian covers of finite type surfaces, and then for the infinite type case, we exhaust by finite type subsurfaces.
The commutator subgroup of $\pi_{1}(S)$ is the kernel of the map $\pi_{1}(S)\to H_{1}(S,\mathbb{Z}),$ so the universal abelian cover could be called the “$\mathbb{Z}$-homology cover” of $S$. We also consider the $\mathbb{Z}/n\mathbb{Z}$-homology cover $\tilde{S}\to S$ corresponding to the kernel of
@@
When $S$ has finite-type, the $\mathbb{Z}/n\mathbb{Z}$-homology cover $\tilde{S}\to S$ is a finite cover, and one can figure out the topological type of $\tilde{S}$ via an Euler characteristic argument and by analyzing how boundary components lift. On the other hand, when $\tilde{S}$ has infinite-type, its $\mathbb{Z}/n\mathbb{Z}$-homology covers all have infinite degree, and it turns out their topology is constrained, just like for $\mathbb{Z}$-homology covers.
@@
Suppose that a cover $\tilde{S}\to S$ corresponds to the subgroup $N<\pi_{1}(S,p)$. We say that the cover is characteristic if $N$ is characteristic, i.e., if $N$ is invariant under every automorphism of $\pi_{1}(S,p)$. The $\mathbb{Z}$ and $\mathbb{Z}/n\mathbb{Z}$-homology covers above are all characteristic. Motivated by Proposition 1.2 and Theorem 1.3, we ask:
@@
An automorphism of $\pi_{1}(S,p)$ is called geometric if it is induced by a homeomorphism $S\to S$ that preserves the basepoint $p$, a subgroup $N<\pi_{1}(S,p)$ is geometrically characteristic if $N$ is invariant under all geometric automorphisms of $\pi_{1}(S,p)$. A cover $\tilde{S}\to S$ is geometrically characteristic if it corresponds to a geometrically characterstic subgroup of $\pi_{1}(S,p)$. Of course, any characteristic cover is geometrically characteristic. Note that geometrically characteristic subgroups are normal, since inner automorphisms of $\pi_{1}(S,p)$ can be realized via point pushing homeomorphisms, and hence geometrically characteristic subgroups are regular. Geometrically characteristic covers were previously considered in [3], where they were defined equivalently as regular covers to which every homeomorphism lifts [^1] There are non-regular covers of closed surfaces where every homeomorphism lifts, see Propositions 6 and 7 in [3]. These are not geometrically characteristic as defined above..
It follows quickly from Theorem 1.5 that any non-universal characteristic cover $\tilde{S}$ of a closed surface is homeomorphic to the Loch Ness monster surface. This was proved earlier by Atarihuana–García–Hidalgo–Quispe–Maluendas [4, Prop 6.1], using the fact that $\tilde{S}$ is also a characteristic cover of a triangle orbifold. However, their argument does not extend to the finite-type setting.
@@
Our interest in Theorem 1.7 is partly inspired by work on big mapping class groups of surfaces. In particular, Aramayona–Leininger–McLeay [3, Proposition 3.3] show that the $\mathbb{Z}/2\mathbb{Z}$ homology cover of the Loch Ness monster surface is the Loch Ness monster surface. They use this to show that the mapping class group $G$ of the once-punctured Loch Ness monster surface admits a continuous injective homomorphism $G\to G$ that is not surjective. We note that there is a gap in the published version of their proof of the covering statement, but they were able to fix it quickly when we notified them about it. Alternatively, their Proposition 3.3 follows from Theorem 1.3, or even from Theorem 1.5 or [4, Proposition 6.1], since the Loch Ness monster surface is itself a characteristic cover of a genus two surface.
Surprisingly, to our knowledge there are not many previous results in which the homeomorphism types of infinite covers of surfaces are studied. In particular, we were surprised that a result like Proposition 1.1 does not currently exist in the literature. However, there are a few results out there that are related to ours. First, a classification of the possible homeomorphism types of infinite-degree regular covers of a closed surface is readily deduced from the fact that the cover and the deck group are quasi-isometric (for instance, see [6, Proposition 5.2]). Then there are older papers of Maskit [14] and Papakyriakopoulos [15] on identifying when certain covers of surfaces are planar; their motivation was some interesting applications to 333-manifolds, including a new proof of the Loop Theorem. There are the papers of Aramayona–Leininger–McLeay [3] and Atarihuana et al [4] referenced above, and there is also a forthcoming paper of Ghaswala–McLeay in which (among other things) the authors construct finite-degree covers from certain maps between potential end spaces of surfaces.
1.1. Ideas of the proofs
@@
Proposition 1.1, in which we show that every noncompact surface covers every surface with non-abelian fundamental group, is very hands-on. Every surface $S$ with non-abelian fundamental group has an essential pair of pants and so is covered by an open pair of pants. The universal abelian cover of an open pair of pants is the Loch Ness monster surface, which is covered by the blooming Cantor tree surface, and every noncompact surface $\tilde{S}$ can be $\pi_{1}$-injectively embedded in the blooming Cantor tree surface, and hence covers the blooming Cantor tree surface. (Here the embedding is as a component of the complement of a set of closed curves and bi-infinite arcs.) Composing, we get a covering map $\tilde{S}\to S$.
Proposition 1.2, in which we identify the universal abelian covers $S^{ab}$ of surfaces $S$, is proved as follows. When $S$ has finite-type, one can identify $S^{ab}$ using the fact that the deck group is free abelian, along with the knowledge of whether the peripheral curves of $S$ lift or not. We deal with infinite-type $S$ by exhausting them with finite-type subsurfaces and saying that $S^{ab}$ is exhausted by the universal abelian covers of the subsurfaces.
@@
Say for simplicity that $S$ is closed. We want to prove that $\tilde{S}$ has one end. Here, $\tilde{S}$ is quasi-isometric to the deck group $Q$. If $Q$ has more than one end, it admits an action on a tree $T$ with finite edge stabilizers and no global fixed point, by Stallings’s Theorem. We then construct a $Q$-equivariant map $\tilde{S}\to T$ that is transverse to the union of edge midpoints, and let $\tilde{C}\subset\tilde{S}$ be the preimage of the set of edge midpoints, which is then a properly embedded 111-submanifold invariant under $Q$. Since edge stabilizers are finite, each component of $\tilde{C}$ is a simple closed curve. Using similar methods and an inductive argument, we can enlarge $\tilde{C}$ until every component of $\tilde{S}\smallsetminus\tilde{C}$ is one-ended.
@@
Theorem 1.7, in which we identify non-locally finite characteristic covers of infinite-type surfaces, is proved by exhausting by finite-type surfaces and using Theorem 1.5, in a way analogous to the infinite-type case of Proposition 1.2.
1.2. Outline of the Paper
In §2 we review the necessary background. We define the surfaces mentioned in the introduction, set conventions, and review actions on trees, ends, Stallings’s Theorem, and the classification of infinite-type surfaces. In §3 we study characteristic covers of surfaces, proving Theorem 1.5, Theorem 1.7, Proposition 1.2 and Theorem 1.3 in subsections 3.1, 3.2, 3.3 and 3.4, respectively. In §4 we prove Proposition 1.1, and then §A is our appendix on ends via group cohomology.
1.3. Acknowledgements
We thank Sumanta Das for point out Goldman’s work and Sébastien Alvarez for several other references related to Proposition 1.1. We also thank the anonymous referee for their comments, which have greatly improved the accuracy and readability of the paper.
@@ 1.3. Acknowledgements
2. Background
2.1. Conventions
All surfaces within this work are assumed to be connected, second countable, orientable, and possibly with boundary. We say the surface is borderless if it has no boundary, and remind the reader that some of the surfaces which appear in our proofs may have non-empty boundary with possibly non-compact boundary components. All borderless surfaces, whether finite or infinite type, are classified by their genus and the topology of their space of ends, see Theorem 2.2. We let $S^{b}_{g,p}$ denote the finite-type surface with $b$ boundary components, genus $g$, and $p$ punctures.
@@ 2.1. Conventions
  3. (3)
    The spotted Loch Ness monster surface is obtained from the Loch Ness monster surface by removing an infinite, discrete, closed set of points.
  4. (4)
@@
2.2. Ends of a topological space
@@
[7, Prop. 8.29] Given two proper geodesic metric spaces $X_{1}$ and $X_{2}$, then any quasi-isometry $f:X_{1}\to X_{2}$ induces a homeomorphism $f_{\mathcal{E}}:\mathcal{E}(X_{1})\to\mathcal{E}(X_{2})$.
2.3. Classification of surfaces
@@
[…]
@@ 2.4. Preliminaries on covers
2.5. Group actions on trees
@@
2.5.1. Stallings’s Theorem and its relative version
@@
Theorem 2.6 is really a result of Swarup [20], but Swarup’s result is stated in terms of group cohomology with group ring coefficients. In Appendix A, we give a gentle introduction to ends via group cohomology and show how the formulation above follows from Swarup’s result.
@@
3. Characteristic covers
In this section we study characteristic covers of surfaces, proving Theorems 1.5, 1.7, Proposition 1.2, and Theorem 1.3 in sections 3.1, 3.2, 3.3, and 3.4, respectively.
3.1. Finite-type surfaces
@@
Let $Q$ be the associated deck group of $\tilde{S}$. First assume that $\tilde{S}$ has no compact boundary components, so $\text{int}(\tilde{S})=\tilde{S}_{\circ}$. As $Q$ is infinite and acts cocompactly on $\tilde{S}$, the genus of $\text{int}(\tilde{S})$ is either 0 or infinity, so the classification of surfaces (Theorem 2.2) implies that it is either the disc or the Loch Ness monster surface. Now, suppose that $\tilde{S}$ has a compact boundary component. As $Q$ is infinite, $\tilde{S}$ has infinitely many compact boundary components. Gluing discs to each compact boundary component, we obtain a surface without boundary, call it $Y$. By Proposition 3.1, $Y$ has a single end, and again the classification of surfaces (Theorem 2.2) implies that $Y$ is either the disc or the Loch Ness monster surface. To finish, observe that $\text{int}(\tilde{S})$ is obtained from $Y$ by puncturing all the infinitely many discs, so then $\text{int}(\tilde{S})$ is either the flute surface or the spotted Loch Ness monster surface. ∎
@@
Let $\pi:\tilde{S}\to S$ be a regular cover of a compact surface $S$ with deck group $Q$. Suppose $Q$ acts on a tree $T$ with finite edge stabilizers, no edge inversions, no global fixed point, and such that the projection to $Q$ of every peripheral element of $\pi_{1}(S)$ acts elliptically on $T$. Then there exists a $Q$-equivariant map $\tilde{f}:\tilde{S}\to T$ such that each component of $\partial\tilde{S}$ maps to a vertex under $\tilde{f}$. Moreover, if $Z$ is the set of edge midpoints in $T$, then the preimage $\tilde{C}:=\tilde{f}^{-1}(Z)$ is a properly embedded 111-submanifold of $\tilde{S}$, each component of which is a simple closed curve, and its projection $C:=\pi(\tilde{C})$ is a compact 111-submanifold of $S$ each component of which is non-peripheral and essential.
Above, there may be components of $C$ that are parallel, which is why we call $C$ a 111-submanifold and not a multicurve.
@@
Figure 1. The map $\tilde{f}_{1}$ from Proposition 3.3 restricted to a triangle in the triangulation $\tilde{\tau}$; the green, blue, and red regions map onto the green, blue, and red segments, respectively, in the tripod; the restriction to any one of the six subtriangles is an orthogonal projection map. The preimage of an edge midpoint $z\in T$ is shown in orange.
@@
First, suppose that there is a component of $C_{1}$ that bounds a disk. Let $\Delta_{1},\ldots,\Delta_{k}$ be the outermost disks bounded by components of $C_{1}$, so that if $c$ is a component of $C_{1}$ bounding a disk, then there exists a unique $j\in\{1,\ldots,k\}$ such that $c\subset\Delta_{j}$. Choose a lift $\tilde{\Delta}_{j}$ in $\tilde{S}$ for each of the $\Delta_{j}$, and for $q\in Q$, let $\tilde{\Delta}_{j}^{q}=q\cdot\tilde{\Delta}_{j}$. Let $e_{j}^{q}$ be the edge with vertices $v_{j}^{q},w_{j}^{q}\in T$ whose midpoint is $\tilde{f_{1}}(\tilde{\partial}\Delta_{j}^{q})$. By the construction of $C_{1}$, all the 111-cells in $\tilde{\tau}$ that intersect ~$\tilde{\partial}\Delta_{j}^{q}$ are mapped to segments in $T$ that all contain the edge $e_{j}^{q}$. If the 0-cells in $\tilde{\Delta}_{j}^{q}$ are mapped into the $v_{j}^{q}$ side of $T\smallsetminus\{e_{j}^{q}\}$, we can define a new map
@@
Since the action of the deck group $Q$ produced by Lemma 3.2 has no global fixed point, the image of the map in Proposition 3.3 has to contain an edge midpoint, the 111-submanifold $C\subset S$ it produces is non-empty. Take $\gamma$ to be any component of $C$. Any component of $\pi^{-1}(\gamma)$ is a component of $\tilde{C}$, and hence is compact. ∎
@@
If $\pi:\tilde{S}\to S$ is a regular cover of a compact surface and $C\subset S$ is a compact 111-submanifold such that every component of $\pi^{-1}(C)\subset\tilde{S}$ is compact. Then there exists a compact 111-submanifold $D\subset S$ containing $C$ such that
@@
3.1.1. Proof of Proposition 3.1
We argue by way of contradiction: assume that $\tilde{S}_{\circ}$ has more than one end. Let $Q$ be the deck group associated to the covering $\pi$. Apply Proposition 3.3 to obtain a tree $T$, a peripherally elliptic action of $Q$ on $T$ with no global fixed point, and the 1-submanifolds $\tilde{C}$ and $C$ of $\tilde{S}$ and $S$, respectively, and apply Lemma 3.5 to get the 111-submanifold $D$ on $S$. Recall that the action of $Q$ on $T$ has finite edge stabilizers, no edge inversions, no global fixed point, and every element of $Q$ that is the image of a peripheral element in $\pi_{1}(S)$ acts elliptically on $T$. The contradiction will arise by showing that the action of $Q$ has a global fixed point.
@@
Figure 2. Standard generators for $\pi_{1}(S)$ for $S$ on a genus two surface with three punctures are given by the elements $a_{1},\ldots,a_{6}$.
Figure 3. A generating set for a thrice-punctured torus consisting of non-separating curves, namely $a_{1},a_{2},a_{3},a_{4}$, where products of pairs of distinct generators are either non-separating, or are freely homotopic to $\delta$. For a similar example on a twice-punctured torus, just delete one of the two boundary components that is not $\delta$.
@@
3.2. Non-locally finite covers of infinite-type surfaces